# Physics-Inspired Extrapolation for efficient error mitigation and hardware certification

### v2

@@ Abstract
Quantum error mitigation is essential for the noisy intermediate-scale quantum era, and will remain relevant for early fault-tolerant quantum computers, where logical error rates are still significant. However, most QEM methods incur an exponential sampling overhead to achieve unbiased estimates, limiting their practical applicability. Recently, error mitigation by restricted evolution was shown to estimate expectation values with constant sampling overhead, albeit with a small bias that grows with circuit size and noise level. Building upon the EMRE framework, here, we propose physics-inspired extrapolation, a linear circuit runtime protocol that achieves enhanced accuracy without incurring substantial overhead. Unlike traditional zero-noise extrapolation methods, PIE provides an operational interpretation of its fitting parameters and converges to unbiased estimates as noise decreases. Distinctively, the slope of the extrapolation fit corresponds to the max-relative entropy between the ideal and noisy circuits, enabling quantitative hardware certification alongside error mitigation, with no additional computational overhead. We also demonstrate the efficacy of this method on IBMQ hardware and apply it to simulate 84-qubit quantum dynamics efficiently. Our results show that PIE yields accurate, low-variance error mitigated estimates, establishing it as a practical and scalable strategy for both error mitigation and hardware certification for near-term and early fault-tolerant quantum computers.
@@ I Introduction
The ubiquitous presence of noise in various quantum hardware modalities presents a fundamental obstacle in realizing an ultimate quantum machine capable of executing industry-relevant applications. Examples of such applications include simulating FeMoco molecules [1], modeling spin defects [2], designing solid-state battery materials [3], optimizing solar cell design [4], and accelerating drug discovery [5], tasks that remain intractable for even the current best classical computers [6, 7].
To overcome the above challenge and achieve fault tolerance in quantum computers, quantum error correction (QEC) [8] emerges as the ultimate tool. QEC refers to a set of techniques that protect quantum information against errors arising due to both environmental effects (such as decoherence) and imperfect gate operations (such as random bit-flip error). Despite the promise of fault-tolerant quantum computing, today’s quantum processors remain constrained to the noisy intermediate-scale quantum (NISQ) regime [9, 10], characterized by a limited number of qubits and susceptibility to various noise sources. Implementing fully error-corrected quantum algorithms [11] capable of outperforming classical computations remains infeasible due to prohibitive resource overheads, with each logical qubit requiring hundreds to thousands of physical qubits. Consequently, until fully fault-tolerant quantum systems are realized, we must shift our focus to alternative strategies [10] for enhancing computational accuracy on existing hardware.
@@
One particularly promising approach is quantum error mitigation (QEM) [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Unlike QEC, QEM techniques do not require the costly overhead of encoding logical qubits. They aim to suppress or eliminate the impact of noise through post-processing of measurement outcomes, enabling more accurate estimation of quantum observables even within the limitations of current NISQ devices. These methods are algorithmic schemes designed to reduce the bias (absolute error between the obtained and ideal results) in the expectation value, and as such do not have a universal or a unique definition. Among the most widely used QEM methods is zero-noise extrapolation (ZNE) [27, 14], where noise is treated as a parameter. By systematically amplifying the noise in the quantum circuit, one can obtain a series of noisy expectation values and extrapolate them back to the zero-noise limit, thereby estimating the ideal result. Another prominent method is the probabilistic error cancellation (PEC) [14], which provides unbiased estimates of the expectation value by using quasi-probabilistic decomposition of the gates in the circuit. However, the unbiased estimate is achieved using a sampling overhead that scales exponentially with the number of gates in the circuit, rendering it impractical for large circuits.
@@
Although current noisy quantum hardware lack enough high-quality physical qubits to construct a sufficient number of logical qubits for algorithmic relevance, an intermediate “early fault-tolerant (EFT)” era is anticipated before full fault-tolerance is achieved. In this EFT era, processors may offer a few hundred to a few thousand logical qubits enabled by imperfect error correction schemes. Noise will not be fully eliminated, and algorithms will still suffer performance losses. Consequently, error mitigation will remain crucial, complementing partial or limited error correction to mitigate against logical noise [28, 29]. This approach is expected to produce the highest possible performance using EFT devices. Thus, error mitigation schemes that do not incur large computational overhead will play a pivotal role in maximizing efficiency and accuracy of the outcomes during the EFT era.
@@
Here, we introduce an error mitigation technique based on the constant runtime Error Mitigation by Restricted Evolution (EMRE) protocol that we recently proposed [26]. The primary contribution of this manuscript is the development and experimental demonstration of the EMRE-inspired extrapolation, termed Physics-Inspired Extrapolation (PIE), on IBMQ devices, with a benchmarking against standard resource-efficient ZNE approaches. We show that the PIE method exhibits four principal advantages (see Fig. 1). First, parameters in the extrapolation function used in PIE have a clear operational interpretation, which is often missing in conventional extrapolation-based mitigation schemes; second, the computational runtime scales linearly with respect to the number of data points chosen for extrapolation, making it computationally efficient; third, the associated estimation variance remains low, ensuring reliability in the results; and fourth, and most distinctively, PIE enables direct quantification of the distance between the ideal and noisy circuits in terms of maximum relative entropy, evaluating noise impact and enabling hardware certification. We demonstrate that in the low-noise regime, PIE method provides accurate estimates of the expectation values using only linear runtime, making it a strong candidate for deployment on EFT quantum processors, where hardware resources and error rates remain significant constraints. Additionally, we also develop another error mitigation protocol, the inverse EMRE, which leverages hardware noise characterization to obtain the generalized quasi-probabilistic decomposition (GQPD) of each gate in the circuit to perform EMRE. Owing to noise characterization and the computation of optimal GQPD, the inverse EMRE method has comparatively longer runtime and larger deviations than the PIE method, and so, we have presented the full discussion of the inverse EMRE method, including its implementation and numerical and experimental results, in the supplementary information.
@@ II Physics-Inspired Extrapolation (PIE)
Characterizing noise in physical hardware is a challenging and resource-intensive task. Since error mitigation techniques aim to estimate the ideal expectation values, procedures that avoid any extra computational cost, such as those incurred from noise characterization, are generally preferred. In this scheme of things, ZNE has proven to be very practical [30]. ZNE works by systematically increasing the circuit fault rate, getting the erroneous expectation value at the different fault rates, and then extrapolating the expectation values to the zero-noise limit. To enhance circuit fault rate to implement ZNE, one of the widely used methods is through circuit folding, where identities are added to the circuit, increasing the length of the circuit and systematically increasing noise [17, 31]. For example, if we assume that the whole circuit is represented by a unitary operator $U$, then folding the circuit once would mean that we implement the circuit $U\circ U^{\dagger}\circ U$ which, in ideal non-erroneous circumstances, should yield the same result as the original case. However, due to the faulty nature of the gates, circuit folding increases the errors in the circuit. By folding the circuit multiple times, we can get erroneous results as a function of the circuit fault rate or the number of circuit foldings. The expectation values obtained in the procedure are then extrapolated to the zero-noise limit.
@@
An uncertainty that ZNE presents is in the choice of the extrapolation function. For example, in [27] a linear extrapolation function was employed to approximate the ideal expectation value, while [14] demonstrated that the ideal expectation value can be approximated using a polynomial function for small circuit fault rates, whereas [32] showed that exponential extrapolation offers improved bias compared to other extrapolation techniques. Other extrapolation functions used for error mitigation can be found in the review article [12]. In addition, for extrapolations going beyond Richardson extrapolation, there are no analytical expressions for bias or variance in most cases [12]. Furthermore, no operational meaning has been given to the extrapolation methods. Intuitive explanations on why the exponential extrapolation method works have been provided, but no solid rigorous explanation/reasoning has been given in regards to which exponential function should be used for extrapolation.
@@
Here, we give a resolution to the above by introducing the error mitigation approach PIE, that is free from heuristic methods. Each step in the protocol, including the choice of extrapolation fit, is guided by operational meaning. Using the extrapolation-based techniques and the EMRE framework, PIE can efficiently estimate the expectation value without requiring explicit knowledge of the underlying hardware noise. We determine an extrapolation function that enables an accurate prediction of the ideal (noiseless) expectation value. Accuracy improves considerably in low-noise regimes, as theoretical approximations converge toward exact values as the probability of noise decreases (see Eq. (4)). We validate the effectiveness of PIE through extensive numerical simulations and experimental demonstrations on IBMQ hardware, showing that it yields reliable, noise-robust estimates of expectation values. We benchmark our results against standard ZNE methods using various extrapolation models, including linear, quadratic, and exponential fits. We find that the expectation value estimate using our PIE method yields highly accurate results while avoiding the large standard deviations often observed with exponential ZNE. Thus, we demonstrate that PIE provides a practically motivated and efficient error mitigation technique, bringing the quantum utility era closer to reality. In the following, we discuss how PIE leverages the EMRE framework and circuit folding to get the extrapolation function and estimate the expectation value efficiently.
@@
EMRE relies on the GQPD of an ideal gate and mitigates errors by restricting the decomposition to its positive component, thereby approximating the ideal operation [26]. To achieve minimum bias, it is necessary to employ the optimal GQPD. However, obtaining this optimal decomposition requires hardware noise characterization, which in itself is both a time-consuming and a resource-intensive process. Even if a (not necessarily optimal) GQPD is known, EMRE can be implemented using only a constant sample overhead. This constant sample overhead comes at the cost of a finite bias which grows exponentially with the circuit size. Since our objective is to mitigate errors without incurring significant runtime or bias, we need to find a way to circumvent noise characterization. In order to do that, we propose constructing a GQPD for each ideal gate (and thus for the ideal circuit) using the noisy gates. We show below that it is always possible to construct a GQPD (not necessarily optimal) of the ideal unitary operation from erroneous ones using maximum relative entropy.
@@
where $\mathcal{M}$ is some quantum channel. Note that if $\mathcal{U}$ denotes an ideal unitary circuit, then $\mathcal{E}$ can be considered as the effective noise channel defined as $\mathcal{E}=\mathcal{F}\circ\mathcal{U}^{\dagger}$ where $\mathcal{F}$ is a quantum channel representing the noisy circuit. Interestingly, PIE offers a hidden advantage by not requiring to compute the GQPD for any gate in the circuit and using their erroneous implementations instead. In standard EMRE, the GQPD of all the gates in the circuit must be known, and the original circuit is replaced with a modified one sampled according to the GQPD associated with each gate. Here, only the gates affected by noise are effectively approximated by their noisy counterparts. As a result, the overall product of the generalized robustness of all gates remains small, as we see from our experimental and numerical results, which further contributes to a reduced bias. However, the associated $s$ remains unknown but we later show that it can be determined using PIE.
@@
Since EMRE works by approximating the ideal unitary operation to the positive part of its GQPD, we can approximate the expectation value of an observable, say $\mathcal{O}$ as follows
@@
where $\langle\mathcal{O}\rangle_{\mathcal{M}}:={\rm Tr}[\mathcal{O}\mathcal{M}(|0\rangle\langle 0|)]$ for any quantum channel, $\mathcal{M}$. Note that the smaller the noise probability, the closer the RHS in the above equation will be to the ideal expectation value in the LHS. Now, to enhance the errors in the circuit, we apply global circuit folding. We assume that the same noise acts throughout the circuit even after increasing the circuit depth, and note that the max relative entropy between $\mathcal{U}^{\dagger}$ and $\mathcal{E}\circ\mathcal{U}^{\dagger}$ is the same as the max relative entropy between $\mathcal{U}$ and $\mathcal{E}\circ\mathcal{U}$. So, after $n$ circuit foldings, we can approximate the expectation value of $\mathcal{O}$ as follows:
@@
where $\lambda=0$ indicates the case without error. Further, taking the log on both the sides, we can cast the above expression as
@@
A notable feature of the extrapolation in Eq.(8) is that the slope of the linear curve corresponds to the logarithm of the maximum relative entropy between the ideal and the noisy circuits, thereby quantifying their operational distance. The max-relative entropy serves as a measure of distinguishability and is widely used in quantum information theory, particularly in tasks such as quantum channel discrimination [33, 34, 35, 36] and resource quantification [37, 38, 39, 40, 41, 42, 43].
@@
In our setting, the max-relative entropy serves as a measure of how far the noisy circuit deviates from its ideal counterpart. In the broader literature, estimating the distance between ideal and noisy channels is known to be a challenging task. Tomography techniques, such as quantum process tomography, allow one to fully characterize the quantum channel but become impractically expensive as system size grows, due to its exponential scaling [44, 45, 46]. Fidelity estimation protocols, such as randomized benchmarking, offer scalable alternatives by estimating average gate performance through statistical overlap between ideal and noisy states [47, 48, 49]. Yet, all these methods typically require some assumptions or knowledge of the noise or ideal channel’s action on the input states in order to compare and reconstruct the process [50]. In contrast, PIE offers a resource efficient approach to estimate this distance via the max relative entropy without requiring prior knowledge of the ideal circuit. This perspective naturally frames PIE not only as an error-mitigation protocol but also as a quantum circuit certification tool that operates in parallel with expectation-value estimation, without incurring additional overhead. Such dual-purpose functionality aligns with emerging trends in quantum system certification, where protocols aim to validate circuit integrity alongside computational output [51].
@@
Figure 3: (Color online) Experimental error-mitigated global magnetization $M_{z}$ on the two Trotter steps simulation of the one-dimensional Ising chain with $N=84$ spins. (a) PIE extrapolation from the noise amplified expectation values for $t=1.5$. (b) Comparison of extrapolation methods. (c) Topology of the IBM Eagle superconducting processor ibm kyiv with the ring of qubits selected for the experiments in dark blue. All experiments were performed using four extrapolation points and 4096 shots.

### v3

@@ Abstract
Quantum error mitigation (QEM) is essential for the noisy intermediate-scale quantum era, and will remain relevant for early fault-tolerant quantum computers, where logical error rates are still significant. However, most QEM methods incur an exponential sampling overhead to achieve unbiased estimates, limiting their practical applicability. Recently, error mitigation by restricted evolution (EMRE) was shown to estimate expectation values with constant sampling overhead, albeit with a small bias that grows with circuit size and noise level. Building upon the EMRE framework, here, we propose physics-inspired extrapolation (PIE), a linear circuit runtime protocol that achieves enhanced accuracy without incurring substantial overhead. Unlike traditional heuristic zero-noise extrapolation methods, PIE provides a theoretical foundation for the operational interpretation of its fitting parameters and demonstrates convergence to unbiased estimates as noise decreases. Distinctively, the slope of the extrapolation fit corresponds to the max-relative entropy between the ideal and noisy circuits, enabling quantitative hardware certification alongside error mitigation, with no additional computational overhead. We also demonstrate the efficacy of this method on IBMQ hardware and apply it to simulate 84-qubit quantum dynamics efficiently. Our results show that PIE yields accurate, low-variance error mitigated estimates, establishing it as a practical and scalable strategy for both error mitigation and hardware certification for near-term and early fault-tolerant quantum computers.
@@ I Introduction
The ubiquitous presence of noise in various quantum hardware modalities presents a fundamental obstacle in realizing an ultimate quantum machine capable of executing industry-relevant applications. Examples of such applications include simulating FeMoco molecules Lee et al. (2021), modeling spin defects Baker et al. (2024), designing solid-state battery materials Delgado et al. (2022), optimizing solar cell design Motlagh et al. (2024), and accelerating drug discovery Santagati et al. (2024), tasks that remain intractable for even the current best classical computers Hoefler et al. (2023); Lee et al. (2023).
To overcome the above challenge and achieve fault tolerance in quantum computers, quantum error correction (QEC) Brun (2013) emerges as the ultimate tool. QEC refers to a set of techniques that protect quantum information against errors arising due to both environmental effects (such as decoherence) and imperfect gate operations (such as random bit-flip error). Despite the promise of fault-tolerant quantum computing, today’s quantum processors remain constrained to the noisy intermediate-scale quantum (NISQ) regime Preskill (2018); Bharti et al. (2022), characterized by a limited number of qubits and susceptibility to various noise sources. Implementing fully error-corrected quantum algorithms Gidney (2025) capable of outperforming classical computations remains infeasible due to prohibitive resource overheads, with each logical qubit requiring hundreds to thousands of physical qubits. Consequently, until fully fault-tolerant quantum systems are realized, we must shift our focus to alternative strategies Bharti et al. (2022) for enhancing computational accuracy on existing hardware.
@@
One particularly promising approach is quantum error mitigation (QEM) Cai et al. (2023); Li and Benjamin (2017a); Temme et al. (2017); Endo et al. (2018a); Bonet-Monroig et al. (2018); Giurgica-Tiron et al. (2020); Sun et al. (2021); Shaw (2021); Lowe et al. (2021b); Koczor (2021); Wang et al. (2021); Suzuki et al. (2022b); Bultrini et al. (2023); Liu et al. (2024); Saxena and Kyaw (2024). Unlike QEC, QEM techniques do not require the costly overhead of encoding logical qubits. They aim to suppress or eliminate the impact of noise through post-processing of measurement outcomes, enabling more accurate estimation of quantum observables even within the limitations of current NISQ devices. These methods are algorithmic schemes designed to reduce the bias (absolute error between the obtained and ideal results) in the expectation value, and as such do not have a universal or a unique definition. Among the most widely used QEM methods is zero-noise extrapolation (ZNE) Li and Benjamin (2017b); Temme et al. (2017), where noise is treated as a parameter. By systematically amplifying the noise in the quantum circuit, one can obtain a series of noisy expectation values and extrapolate them back to the zero-noise limit, thereby estimating the ideal result. Another prominent method is the probabilistic error cancellation (PEC) Temme et al. (2017), which provides unbiased estimates of the expectation value by using quasi-probabilistic decomposition of the gates in the circuit. However, the unbiased estimate is achieved using a sampling overhead that scales exponentially with the number of gates in the circuit, rendering it impractical for large circuits.
@@
Although current noisy quantum hardware lack enough high-quality physical qubits to construct a sufficient number of logical qubits for algorithmic relevance, an intermediate “early fault-tolerant (EFT)” era is anticipated before full fault-tolerance is achieved. In this EFT era, processors may offer a few hundred to a few thousand logical qubits enabled by imperfect error correction schemes. Noise will not be fully eliminated, and algorithms will still suffer performance losses. Consequently, error mitigation will remain crucial, complementing partial or limited error correction to mitigate against logical noise Suzuki et al. (2022a); Dutkiewicz et al. (2025). This approach is expected to produce the highest possible performance using EFT devices. Thus, error mitigation schemes that do not incur large computational overhead will play a pivotal role in maximizing efficiency and accuracy of the outcomes during the EFT era.
@@
Here, we introduce an error mitigation technique based on the constant runtime Error Mitigation by Restricted Evolution (EMRE) protocol that we recently proposed Saxena and Kyaw (2024). The primary contribution of this manuscript is the development and experimental demonstration of the EMRE-inspired extrapolation, termed Physics-Inspired Extrapolation (PIE), on IBMQ devices, with a benchmarking against standard resource-efficient ZNE approaches. While the protocol is mathematically equivalent to a specific choice of a ZNE method, we show that the PIE technique exhibits four principal advantages (see Fig. 1). First, parameters in the extrapolation function used in PIE have a clear operational interpretation, which is often missing in conventional extrapolation-based mitigation schemes; second, the computational runtime scales linearly with respect to the number of data points chosen for extrapolation, making it computationally efficient; third, the associated estimation variance remains low, ensuring reliability in the results; and fourth, and most distinctively, PIE enables direct quantification of the distance between the ideal and noisy circuits in terms of maximum relative entropy, evaluating noise impact and enabling hardware certification. We demonstrate that in the low-noise regime, PIE method provides accurate estimates of the expectation values using only linear runtime, making it a strong candidate for deployment on EFT quantum processors, where hardware resources and error rates remain significant constraints. Additionally, we also develop another error mitigation protocol, the inverse EMRE, which leverages hardware noise characterization to obtain the generalized quasi-probabilistic decomposition (GQPD) of each gate in the circuit to perform EMRE. Owing to noise characterization and the computation of optimal GQPD, the inverse EMRE method has comparatively longer runtime and larger deviations than the PIE method, and so, we have presented the full discussion of the inverse EMRE method, including its implementation and numerical and experimental results, in the supplementary information.
@@ II Physics-Inspired Extrapolation (PIE)
Characterizing noise in physical hardware is a challenging and resource-intensive task. Since error mitigation techniques aim to estimate the ideal expectation values, procedures that avoid any extra computational cost, such as those incurred from noise characterization, are generally preferred. In this scheme of things, ZNE has proven to be very practical Kim et al. (2023). ZNE works by systematically increasing the circuit fault rate, getting the erroneous expectation value at the different fault rates, and then extrapolating the expectation values to the zero-noise limit. To enhance circuit fault rate to implement ZNE, one of the widely used methods is through circuit folding, where identities are added to the circuit, increasing the length of the circuit and systematically increasing noise Giurgica-Tiron et al. (2020); Majumdar et al. (2023). For example, if we assume that the whole circuit is represented by a unitary operator $U$, then folding the circuit once would mean that we implement the circuit $U\circ U^{\dagger}\circ U$ which, in ideal non-erroneous circumstances, should yield the same result as the original case. However, due to the faulty nature of the gates, circuit folding increases the errors in the circuit. By folding the circuit multiple times, we can get erroneous results as a function of the circuit fault rate or the number of circuit foldings. The expectation values obtained in the procedure are then extrapolated to the zero-noise limit.
@@
An uncertainty that ZNE presents is in the choice of the extrapolation function. For example, in Li and Benjamin (2017b) a linear extrapolation function was employed to approximate the ideal expectation value, while Temme et al. (2017) demonstrated that the ideal expectation value can be approximated using a polynomial function for small circuit fault rates, whereas Endo et al. (2018b) showed that exponential extrapolation offers improved bias compared to other extrapolation techniques. Other extrapolation functions used for error mitigation can be found in the review article Cai et al. (2023). In addition, for extrapolations going beyond Richardson extrapolation, there are no analytical expressions for bias or variance in most cases Cai et al. (2023). Furthermore, no operational meaning has been given to the extrapolation methods. Intuitive explanations on why the exponential extrapolation method works have been provided, but no solid rigorous explanation/reasoning has been given regarding which exponential function should be used for extrapolation.
@@
Here, we give a resolution to the above by introducing the error mitigation approach PIE, which is free from heuristic methods. Each step in the protocol, including the choice of extrapolation fit, is guided by operational meaning. Using the extrapolation-based techniques and the EMRE framework, PIE can efficiently estimate the expectation value without requiring explicit knowledge of the underlying hardware noise. We determine an extrapolation function that enables an accurate prediction of the ideal (noiseless) expectation value. Accuracy improves considerably in low-noise regimes, as theoretical approximations converge toward exact values as the probability of noise decreases (see Eq. (4)). We validate the effectiveness of PIE through extensive numerical simulations and experimental demonstrations on IBMQ hardware, showing that it yields reliable, noise-robust estimates of expectation values. We benchmark our results against standard ZNE methods using various extrapolation models, including linear, quadratic, and exponential fits. We find that the expectation value estimate using our PIE method yields highly accurate results while avoiding the large standard deviations often observed with exponential ZNE. Thus, we demonstrate that PIE provides a practically motivated and efficient error mitigation technique, bringing the quantum utility era closer to reality. Importantly, our results do not contradict the fundamental limitations established in Quek et al. (2024). At high noise probabilities, the estimates from PIE begin to deviate from the ideal values as PIE does not address the underlying source of exponential bias arising from the omitted terms in EMRE. However, as gate fidelities continue to improve, PIE becomes increasingly effective, offering a practical and efficient error mitigation strategy that uses a non-heuristic and practically motivated extrapolation fit whose parameters have an underlying operational interpretation. In the following, we discuss how PIE leverages the EMRE framework and circuit folding to get the extrapolation function and estimate the expectation value efficiently.
@@
EMRE relies on the GQPD of an ideal gate and mitigates errors by restricting the decomposition to its positive component, thereby approximating the ideal operation Saxena and Kyaw (2024). To achieve minimum bias, it is necessary to employ the optimal GQPD. However, obtaining this optimal decomposition requires hardware noise characterization, which in itself is both a time-consuming and a resource-intensive process. Even if a (not necessarily optimal) GQPD is known, EMRE can be implemented using only a constant sample overhead. This constant sample overhead comes at the cost of a finite bias which grows exponentially with the circuit size. Since our objective is to mitigate errors without incurring significant runtime or bias, we need to find a way to circumvent noise characterization. In order to do that, we propose constructing a GQPD for each ideal gate (and thus for the ideal circuit) using the noisy gates. We show below that it is always possible to construct a GQPD (not necessarily optimal) of the ideal unitary operation from erroneous ones using maximum relative entropy.
@@
where $\mathcal{M}$ is some quantum channel. Note that if $\mathcal{U}$ denotes an ideal unitary circuit, then $\mathcal{E}$ can be considered as the effective noise channel defined as $\mathcal{E}=\mathcal{F}\circ\mathcal{U}^{\dagger}$ where $\mathcal{F}$ is a quantum channel representing the noisy circuit. Interestingly, PIE offers a hidden advantage by not requiring the computation of the GQPD for any gate in the circuit and using their erroneous implementations instead. In standard EMRE, the GQPD of all the gates in the circuit must be known, and the original circuit is replaced with a modified one sampled according to the GQPD associated with each gate. Here, only the gates affected by noise are effectively approximated by their noisy counterparts. As a result, the overall product of the generalized robustness of all gates remains small, as we see from our experimental and numerical results, which further contributes to a reduced bias. However, the associated $s$ remains unknown, but we later show that it can be determined using PIE.
@@
Since EMRE works by approximating the ideal unitary operation to the positive part of its GQPD, we can approximate the expectation value of an observable, say $\mathcal{O}$, as follows
@@
where $\langle\mathcal{O}\rangle_{\mathcal{M}}:={\rm Tr}[\mathcal{O}\mathcal{M}(|0\rangle\langle 0|)]$ for any quantum channel, $\mathcal{M}$. Note that the smaller the noise probability, the closer the RHS in the above equation will be to the ideal expectation value in the LHS. As a remark, we also emphasize that from the above approximation relation, one might incorrectly infer that $s$ depends on the choice of observable. As we show below, for any nontrivial observable $\mathcal{O}$, the slope of the extrapolation fit (which depends solely on $s$) remains invariant as long as the circuit and noise are fixed, implying that $s$ is independent of the observable. Indeed, since $s$ is a distance measure, it must remain independent of the observable and depend only on the ideal and noisy operations, as defined in Eq. 2.
Now, to enhance the errors in the circuit, we apply global circuit folding. We assume that the same noise acts throughout the circuit even after increasing the circuit depth, and note that the max relative entropy between $\mathcal{U}^{\dagger}$ and $\mathcal{E}\circ\mathcal{U}^{\dagger}$ is the same as the max relative entropy between $\mathcal{U}$ and $\mathcal{E}\circ\mathcal{U}$. So, after $n$ circuit foldings, we can approximate the expectation value of $\mathcal{O}$ as follows:
@@
where $\lambda=0$ indicates the case without error. Further, taking the log on both sides, we can cast the above expression as
@@
We also emphasize that PIE can, in principle, address SPAM errors, though doing so introduces an additional approximation error if no complementary mitigation methods are employed. SPAM errors can be separated into two components: incorrect state preparation and incorrect measurement. We can view state preparation as part of the quantum circuit itself. In other words, if we only require that the qubits are initialized in the $|0\rangle^{\otimes n}$ state perfectly, then state preparation may be regarded as a pre-computation step whose imperfections simply contribute to the overall gate-level noise. For measurement errors, one way is to use alternate/complimentary EM method such as T-Rex which is widely used in practice and has worked well. Alternatively, if no such complimentary EM method is used, one may decompose the ideal observable (similar to GQPD) into a linear combination of implementable observable/s (which will form the positive component of the decomposition) and an additional matrix (the negative part), in line with what we did for other operations in the circuit (see for reference, Eq. (3)). Next, we can approximate the ideal observable $\mathcal{O}$ to be the erroneous observable $\mathcal{O}^{\rm err}$ scaled by a factor, say $s_{m}$, i.e., approximate $\mathcal{O}$ as $\mathcal{O}\approx s_{m}\mathcal{O}^{\rm err}$. Since folding cannot be applied to this observable, $s_{m}$ cannot be absorbed into the max-relative entropy of the circuit $s$. Consequently, we will have an extra parameter in Eqs. (4) and (5), and the ideal outcome after $n$ circuit folds can be approximated as $\langle\mathcal{O}\rangle_{\rm ideal}\approx s_{m}s^{(2n+1)}\langle\mathcal{O}\rangle_{\rm err}\,.$ Taking logarithms introduces an additional parameter, which must be learned. This parameter can be estimated using simple observables within a learning-based procedure and subsequently applied to the noisy measurements of the target observable to recover its ideal expectation value.
@@
A notable feature of the extrapolation in Eq.(16) is that the slope of the linear curve corresponds to the logarithm of the maximum relative entropy between the ideal and the noisy circuits, thereby quantifying their operational distance. The max-relative entropy serves as a measure of distinguishability and is widely used in quantum information theory, particularly in tasks such as quantum channel discrimination Wang and Wilde (2019); Li et al. (2022); Koßmann and Wilde (2024); Fang et al. (2025) and resource quantification Datta (2009); Bu et al. (2017); Gour and Winter (2019); Saxena et al. (2020); Wang et al. (2020); Saxena and Gour (2022); Chitambar and Gour (2019).
@@
In our setting, the max-relative entropy serves as a measure of how far the noisy circuit deviates from its ideal counterpart. In the broader literature, estimating the distance between ideal and noisy channels is known to be a challenging task. Tomography techniques, such as quantum process tomography, allow one to fully characterize the quantum channel but become impractically expensive as system size grows, due to its exponential scaling Poyatos et al. (1997); Mohseni et al. (2008); Torlai et al. (2023). Fidelity estimation protocols, such as randomized benchmarking, offer scalable alternatives by estimating average gate performance through statistical overlap between ideal and noisy states Knill et al. (2008); Dankert et al. (2009); Helsen et al. (2022). Yet, all these methods typically require some assumptions or knowledge of the noise or ideal channel’s action on the input states in order to compare and reconstruct the process Eisert et al. (2020). In contrast, PIE offers a resource-efficient approach to estimate this distance via the max relative entropy without requiring prior knowledge of the ideal circuit. This perspective naturally frames PIE not only as an error-mitigation protocol but also as a quantum circuit certification tool that operates in parallel with expectation-value estimation, without incurring additional overhead. Such dual-purpose functionality aligns with emerging trends in quantum system certification, where protocols aim to validate circuit integrity alongside computational output Kliesch and Roth (2021).
@@
Figure 3: (Color online) Experimental error-mitigated global magnetization $M_{z}$ on the two Trotter steps simulation of the one-dimensional Ising chain with $N=84$ spins. (a) PIE extrapolation from the noise amplified expectation values for $t=1.5$. The noise-free, experimental unmitigated, and PIE mitigated results are $0.526\pm 0.013$, $0.405\pm 0,002$, and $0.515\pm 0.012$, respectively. (b) Comparison of extrapolation methods. (c) Topology of the IBM Eagle superconducting processor ibm kyiv with the ring of qubits selected for the experiments in dark blue. All experiments were performed using four extrapolation points and 4096 shots.